The earth has a net negative charge of about -4.3 x 105 C. Since this charge is distributed more or less uniformly over the spherical earth, the force due to this charge is the same as if it were all placed at the Earths center. A 6.0 kg satellite is orbiting with an orbital radius of 55000 km. If the satellite were also charged, then there would be an electric force in addition to the gravitational force a. What charge (including the sign) would need to be on the satellite to completely cancel the earths gravitational force on it? b. With this same charge and these same assumptions, what would be the net force on the satellite if instead of being in space it were near the surface of the earth? Hint: Note that the gravitational force is not just mg, you must use Newtons law of gravity since the satellite in part a is not near the earths surface. For part b, you need not do an extensive calculation. The two kinds of force have very similar formulas. Look carefully at their formulas and think about how they vary with distance.
Answer #1

Given is:-

charge on the Earth Q=-4.3 × 10°C

mass of the satellite  m = 6.3 kg

radius  r = 55000 × 103m = 55 × 106771

mass of the Earth  M5.972 x 1024kq

Now,

Part-a

Since the gravitational force is of attractive type therefore the charge on the satellite must be of negative type so that it can counter gravitational force

so

F_e = F_g

rac{KQq}{r^2}= rac{GMm}{r^2}

by plugging all the values we get

(9 × 10°)(-4.3 x 10°)q (55 x 106)2 (6.67× 10-11 )(5.972x 10 24 )(6) (55 × 106 )2

which gives us

0.62C

Part-b

The gravitational force near the Earth's surface is

F,-mag = 6 × 9.8 = 58.8N

and the eletrostatic force will be

KQ where R is the radius of the Earth

by plugging all the values we get

(9x109) (4. 3 x 105) 6300 x 103)2 Fe-(0.62) ×

which gives us

F, 60.45N

Thus the net force on to the sattelite is

F_{Net} = F_e - F_g

FNet 60.45-58.8

FNet 1.65Naway from the earth